# Rapid formation of a very massive star $>\!\!50000\,\mathrm{M}_{\odot}$ and subsequently an IMBH from runaway collisions

### v1

@@ 2.1 NBODY6++GPU
The dense star cluster models are evolved using the state-of-the-art direct force integration code Nbody6++GPU, which is optimised for parallelisation using simultaneously three different levels, bottom-end of many core GPU accelerated parallel computing (Nitadori & Aarseth, 2012; Wang et al., 2015), mid-level OpenMP thread based parallelisation, and upper level MPI parallelisation (Spurzem, 1999). It yields excellent sustained performance on current hybrid massively parallel supercomputers (cf. Spurzem & Kamlah, 2023, and Sect. Hardware settings). It is a successor to the many direct force integration $N$-body codes of gravitational $N$-body problems, which were originally written by Sverre Aarseth (Aarseth, 1985, 1999b, 1999a, 2003b; Aarseth et al., 2008, and sources therein). The recent review by Spurzem & Kamlah (2023) provides a comprehensive overview over the entire research field of collisional dynamics and how Nbody6++GPU fits within this field. That review also provides some key informations about codes, such as PeTaR (Wang et al., 2020a) and BiFROST (Rantala et al., 2023).
@@
The full scale parallel code still retains previously implemented features that are essential for efficient and accurate $N$-body integration on all scales. This includes the Hermite scheme with hierarchical block time-steps (McMillan, 1986; Hut et al., 1995; Makino, 1991; Makino & Aarseth, 1992). The Ahmad-Cohen (AC) neighbour scheme (Ahmad & Cohen, 1973) permits for every star to divide the gravitational forces acting on it into the regular component, originating from distant stars, and an irregular part, originating from nearby stars (”neighbours”). While the AC scheme is very efficient to reduce the number of floating operations necessary, it somewhat disturbs parallelization and acceleration, the code requires both CPU and GPU (Wang et al., 2015). The code also retains Kustaanheimo-Stiefel (KS) regularisation (Stiefel & Kustaanheimo, 1965), which allows fast and accurate integration of hard binaries and few-body systems (Mikkola & Tanikawa, 1999b, a; Mikkola & Aarseth, 1998) as well as close encounters.
@@
Post-Newtonian dynamics of relativistic binaries or fast and close hyperbolic encounters is currently still using an orbit-averaged algorithm, for bound systems (Peters & Mathews, 1963; Peters, 1964); it has been used by different simulations already (Di Carlo et al., 2019, 2020a, 2020b, 2021; Rizzuto et al., 2021, 2022; Arca-Sedda et al., 2021), recently also including spin-dependent kicks at mergers (Arca Sedda et al., 2023a, 2024a, 2024b). For unbound systems Turner (1977) (see also more recent Bae et al., 2017) the relativistic treatment of hyperbolic encounters is still experimental and not yet public in Nbody6++GPU. However, the use of fully resolved relativistic Post-Newtonian (PN) dynamics inside Nbody6 and Nbody6++ has been already pioneered by Kupi et al. (2006), and extended by Aarseth (2012); Brem et al. (2013). A recent paper Sobolenko et al. (2021) shows the current best implementation using PN terms up to order PN3.5 and including spin-spin and spin-orbit interactions. These features also are not yet incorporated in our public code version.
@@
Recently, the 19 Dragon-II simulations, which are formally the successor simulations to the Dragon-I simulations by Wang et al. (2016), executed with Nbody6++GPU have provided pioneering insights into the evolution and dynamics of young massive star clusters found in the Local Universe as well as binary evolution, IMBH growth and gravitational wave sources within them and in the surrounding field (Arca Sedda et al., 2023a, 2024a, 2024b). One key update in these simulations next to the stellar evolution published in Kamlah et al. (2022a), has been the inclusion of general relativistic merger recoil kicks (Arca Sedda et al., 2023b). We are using this code version with important updates about stellar collisions, explained in Appendix C.
@@ 2.2 MOCCA
The fast formation of a VMS through runway collisions of MS stars in dense star clusters has also been demonstrated using star cluster simulation codes based on Hénon (1971) Monte Carlo method (Hénon, 1972a, b, 1975; Stodołkiewicz, 1982, 1986; Giersz, 1998; Joshi et al., 2000; Giersz, 2001, 2006; Giersz et al., 2008; Hypki & Giersz, 2013; Rodriguez et al., 2022; Giersz et al., 2024) for treating stellar dynamics. Gürkan et al. (2004) found that Monte Carlo simulations of star clusters with higher central concentration and short initial relaxation times can undergo core collapse on timescales shorter than the MS lifetime of the most massive stars in the cluster. They found that this early core collapse can lead to an increased collision rate which is likely to result in runaway collisions between MS stars leading to the formation of a VMS. Freitag et al. (2006c, a) carried out Monte Carlo simulations of about a 100 star cluster models with varying initial size ($\rm r_{h}$ ranging between $0.03-5\mathrm{pc}$) and mass ($\rm N_{*}=10^{5}-10^{8}$). These simulations included prescriptions to allow direct collisions between stars, considering both a sticky-sphere assumption for mergers with no mass loss and fractional mass loss of up to 10%. They found that in their densest cluster models that had core collapse times less than about 3 Myr, runaway mergers between MS stars always led to the formation of a VMS that had mass ranging between $\rm 400-4000\ M_{\odot}$. The studies among others have posited that the VMS that forms in this way could be a potentially evolve into an IMBH (Freitag et al., 2007; Freitag, 2008; Goswami et al., 2012; González Prieto et al., 2024).
@@
Simulations of the evolution star cluster models carried out using the Monte Carlo MOCCA code (Hypki & Giersz, 2013; Giersz et al., 2013) have shown that massive clusters with high central densities ($\rm\rho_{c}\gtrsim 10^{7}\ \mathrm{M}_{\odot}\ pc^{-3}$) are likely to form a VMS through the runaway collisions of MS stars (Giersz et al., 2015; Askar et al., 2017a; Hong et al., 2020; Maliszewski et al., 2022; Askar et al., 2023). This process and its potential role in the formation of an IMBH has been described as the FAST scenario in Giersz et al. (2015). In order to trigger this pathway for VMS formation and subsequently IMBH formation, mass segregation of the most massive stars in the cluster must occur on a timescale shorter than their lifetime which is about 3 Myr (Gürkan et al., 2004). High central densities and large sizes of MS stars increases their likelihood to collide and form a VMS. From thousands of star cluster models simulated as part of the MOCCA-Survey Database I (Askar et al., 2017b), it has been shown that the FAST scenario for IMBH formation always occurs when the initial central density exceeds $\rm\rho_{c}\gtrsim 10^{7}\ \mathrm{M}_{\odot}\ pc^{-3}$ (see Fig. 1 in Hong et al., 2020). The mass of the VMS that forms through this pathway is of the order $10^{3}-10^{4}\ \mathrm{M}_{\odot}$ and depends on the total cluster mass and initial central density. Although the formation of a VMS and subsequently an IMBH via the FAST scenario has been extensively explored in previous MOCCA simulations of massive and dense star clusters, these results had not been directly verified with N-body simulations for star clusters with extremely high central densities until now.
@@
In addition to employing the Hénon (1971) Monte Carlo method to treat two-body relaxation, MOCCA utilizes the fewbody code, a direct N-body integrator for small-N gravitational dynamics, to determine the outcomes of close binary-single and binary-binary interactions (Fregeau et al., 2004). For physical collisions between two MS stars in MOCCA, a sticky-sphere assumption with no mass loss is applied. Similar to NBODY6++GPU, the MOCCA code models stellar and binary evolution using prescriptions from the SSE/BSE rapid population synthesis codes (Hurley et al., 2000a, 2002b), which have been subsequently improved and extended with more recent developments (see Kamlah et al., 2022a, and references therein). While the MOCCA code can model a galactic tidal field using a point-mass approximation, no external tidal field was applied in the simulations carried out in this study, in order to remain consistent with the direct N-body run. A detailed description of the latest MOCCA features is provided in Hypki et al. (2022). For this work, MOCCA was updated to incorporate the improved treatment of stellar radii and the rejuvenation process for a VMS that have been described in Appendices A, B and C. Section 3.2 outlines the initial conditions and code modifications applied to MOCCA, while Section 4.3 presents the results on VMS growth and IMBH formation in the MOCCA simulation.
@@ 3.1 Star cluster parameters
Half-mass radius ($r_{\rm hm}$) $0.10\leavevmode\nobreak\ \mathrm{pc}$
Half-mass average density ($\rho_{\rm hm}$) $6.99\cdot 10^{7}\leavevmode\nobreak\ \rm M_{\odot}/{\rm pc}^{3}$
Density profile King ($W_{0}$) $6$ (1)
@@
IMF, min/max stellar mass $0.08-150\leavevmode\nobreak\ \rm M_{\odot}$ (2)
Initial cluster mass ($M_{\rm i}$) $5.86\cdot 10^{5}\leavevmode\nobreak\ \rm M_{\odot}$
@@
In this paper we present one selected star cluster model, all relevant global parameters are given in Table 1. The dynamical evolution of this system is followed by Nbody6++GPU as well as MOCCA, for the purpose of comparison. The initial model was generated by Mcluster (Kuepper et al., 2011; Leveque et al., 2021). The initial number of stars is set to $10^{6}$, with no initial (primordial) binaries. While such binaries in principle can be included (see e.g. Wang et al., 2016; Rizzuto et al., 2021) they are computationally very expensive and also many free parameters for their initial configuration are added. We postpone models with many primordial binaries and a more extended parameter study of initially highly compact star cluster models to future work; in our simulations, however, binaries form and evolve dynamically.The cluster density is modelled via a King (1962) profile with central dimensionless well $W_{0}=6$. Initial stellar masses are distributed according to a Kroupa (2001) initial mass function limited between $0.08-150\leavevmode\nobreak\ \rm M_{\odot}$. All details about our cluster model are summarised in Table 1. The cluster model is initially in virial equilibrium, with no initial mass-segregation, rotation , or any level of fractality (Goodwin & Whitworth, 2004). We do not use an external tidal field i.e., an external gravitational field from the host galaxy, since we are more interested in the dynamical evolution near and inside the half-mass radius. Still, it is possible that stars escape, if their total energy is positive and their distance from the cluster center is larger than 20 initial half-mass radii Aarseth (2003a).
Crossing time and relaxation time taken at the half-mass radius of a star cluster ($t_{\rm cr,hm}$ $t_{\rm rx,hm}$) are important parameters determining the general evolution of a star cluster (Binney & Tremaine, 1987, 2008). The crossing time corresponds to the time in which a star with a typical velocity travels through the cluster under the assumption of virial equilibrium. The half-mass relaxation time corresponds to the time over which the cumulative effect of stellar encounters on the velocity of a star becomes comparable to the star’s initial velocity. In our star cluster model the initial relaxation time is 7.017.017.01\,7.01 Myr, and the initial crossing time is 5.9 $5.9\cdot\ 10^{-4}\,$ Myr.
The mass segregation timescale is important to describe the initial segregation of our star cluster model (Kroupa, 1995). The mass segregation timescale is defined as:
@@
where $\overline{m}$ is the initial average mass of the stars in the cluster, while $m_{\rm max}$ is the largest star mass in the cluster, which are 0.58 $\leavevmode\nobreak\ \rm M_{\odot}$ and 148.49 $\leavevmode\nobreak\ \rm M_{\odot}$, respectively. This, implies a segregation time of $t_{\rm ms}=0.03$ Myr. Moreover, given the large density of the cluster model, this means that encounters between large masses occurs relatively early in the star cluster dynamical history.
[…]
@@
where the effective cross-section is given by $\Sigma_{0}=16\sqrt{\pi}R_{*}^{2}(1+\Theta)$, the Safronov number is defined as $\Theta=9.54\left(M_{*}R_{\odot}/\leavevmode\nobreak\ \rm M_{\odot}R_{*}\right)\left(100\leavevmode\nobreak\ \rm{km\,s^{-1}}/\sigma\right)^{2}$, and $n$ is the numeric density. Vergara et al. (2023) found that the fraction of the total mass in the central object dramatically increase for $t_{\rm coll}<\tau$, being up to 50% for the most extreme runs.
Fig. 1, shows the initial mass of the cluster against the half-mass radius and includes a color bar with the average logarithm of the number of stars. It also displays the relaxation time and the collision time (or critical mass), when both are equals to a time evolution (i.e. age) of $\tau=10\,$ Gyr, which we adopt here as an upper limit for the evolutionary timescales in the simulations we are comparing with. We compare the respective simulations for several grids of direct $N$-body shown as circle symbols in plot representing initial conditions from P+99: Portegies Zwart et al. (1999b), F+13: Fujii & Portegies Zwart (2013), K+15: Katz et al. (2015), W+15: Wang et al. (2015), M+16: Mapelli (2016), S+17: Sakurai et al. (2017), R+18: Reinoso et al. (2018), P+19: Panamarev et al. (2019), R+20: Reinoso et al. (2020), D+20: Di Carlo et al. (2020a), R+20: Rastello et al. (2021), R+21: Rizzuto et al. (2021), B+21: Banerjee (2021), G+21: Gieles et al. (2021), V+21: Vergara et al. (2021), K+22: Kamlah et al. (2022b), R+23: Rizzuto et al. (2023), V+23: Vergara et al. (2023), and AS+23: Arca Sedda et al. (2023b, 2024a, 2024b), Monte-Carlo shown as square symbols in plot representing initial conditions from A+17: Askar et al. (2017b), R+19: Rodriguez et al. (2019), K+20: Kremer et al. (2020b), M+22: Maliszewski et al. (2022), and R+22: Rodriguez et al. (2022) and hybrid simulations diamond symbols in plot representing initial conditions from W+21: Wang et al. (2021), W+22: Wang et al. (2022b), and W24: Wang et al. (2024). The latter are classified as hybrid $N$-body simulations, because they combine particle-particle (e.g. direct $N$-body) and particle-tree (e.g. Barnes-Hut tree) methods to speed-up calculations, and exploit the advantage of a highly parallelised implementation using the code PeTaR (Wang et al., 2020a).

### v2

@@ 2.1 NBODY6++GPU
The dense star cluster models are evolved using the state-of-the-art direct force integration code Nbody6++GPU, which is optimised for parallelisation using simultaneously three different levels, bottom-end of many core GPU accelerated parallel computing (Nitadori2012; Wang2015), mid-level OpenMP thread based parallelisation, and upper level MPI parallelisation (Spurzem1999). It yields excellent sustained performance on current hybrid massively parallel supercomputers (cf. Spurzem2023, and Sect. F). It is a successor to the many direct force integration $N$-body codes of gravitational $N$-body problems, which were originally written by Sverre Aarseth (Aarseth1985; Aarseth1999; Aarseth1999a; Aarseth2003; Aarseth2008, and sources therein). The recent review by Spurzem2023 provides a comprehensive overview over the entire research field of collisional dynamics and how Nbody6++GPU fits within this field. That review also provides some key informations about codes, such as PeTaR (Wang2020a) and BiFROST (Rantala2023).
@@
The full scale parallel code still retains previously implemented features that are essential for efficient and accurate $N$-body integration on all scales. This includes the Hermite scheme with hierarchical block time-steps (McMillan1986; Hut1995; Makino1991; Makino1992). The Ahmad-Cohen (AC) neighbour scheme (Ahmad1973) permits for every star to divide the gravitational forces acting on it into the regular component, originating from distant stars, and an irregular part, originating from nearby stars (”neighbours”). While the AC scheme is very efficient to reduce the number of floating operations necessary, it somewhat disturbs parallelization and acceleration, the code requires both CPU and GPU (Wang2015). The code also retains Kustaanheimo-Stiefel (KS) regularisation (Stiefel1965), which allows fast and accurate integration of hard binaries and few-body systems (Mikkola1999; Mikkola1999a; MikkolaAarseth1998) as well as close encounters.
@@
Post-Newtonian dynamics of relativistic binaries or fast and close hyperbolic encounters is currently still using an orbit-averaged algorithm, for bound systems (Peters1963; Peters1964); it has been used by different simulations already (DiCarlo2019; DiCarlo2020; DiCarlo2020a; DiCarlo2021; Rizzuto2021; Rizzuto2022; ArcaSedda2021a), recently also including spin-dependent kicks at mergers (ArcaSedda2023; ArcaSedda2024; ArcaSedda2024a). For unbound systems Turner1977 (see also more recent Bae2017) the relativistic treatment of hyperbolic encounters is still experimental and not yet public in Nbody6++GPU. However, the use of fully resolved relativistic Post-Newtonian (PN) dynamics inside Nbody6 and Nbody6++ has been already pioneered by Kupi2006, and extended by Aarseth2012; Brem2013. A recent paper Sobolenko2021 shows the current best implementation using PN terms up to order PN3.5 and including spin-spin and spin-orbit interactions. These features also are not yet incorporated in our public code version.
@@
Recently, the 19 Dragon-II simulations, which are formally the successor simulations to the Dragon-I simulations by Wang2016, executed with Nbody6++GPU have provided pioneering insights into the evolution and dynamics of young massive star clusters found in the Local Universe as well as binary evolution, IMBH growth and gravitational wave sources within them and in the surrounding field (ArcaSedda2023; ArcaSedda2024; ArcaSedda2024a). One key update in these simulations next to the stellar evolution published in Kamlah2022, has been the inclusion of general relativistic merger recoil kicks (ArcaSedda2023a). We are using this code version with important updates about stellar collisions, explained in Appendix C.
@@ 2.2 MOCCA
The fast formation of a VMS through runway collisions of MS stars in dense star clusters has also been demonstrated using star cluster simulation codes based on Henon1971 Monte Carlo method (Henon1972a; Henon1972b; Henon1975; Stodolkiewicz1982; Stodolkiewicz1986; Giersz1998; Joshi2000; Giersz2001; Giersz2006; Giersz2008; Hypki2013; Rodriguez2022; Giersz2025) for treating stellar dynamics. Gurkan2004 found that Monte Carlo simulations of star clusters with higher central concentration and short initial relaxation times can undergo core collapse on timescales shorter than the MS lifetime of the most massive stars in the cluster. They found that this early core collapse can lead to an increased collision rate which is likely to result in runaway collisions between MS stars leading to the formation of a VMS. Freitag2006a; Freitag2006 carried out Monte Carlo simulations of about a 100 star cluster models with varying initial size ($\rm r_{h}$ ranging between $0.03-5\mathrm{pc}$) and mass ($\rm N_{*}=10^{5}-10^{8}$). These simulations included prescriptions to allow direct collisions between stars, considering both a sticky-sphere assumption for mergers with no mass loss and fractional mass loss of up to 10%. They found that in their densest cluster models that had core collapse times less than about 3 Myr, runaway mergers between MS stars always led to the formation of a VMS that had mass ranging between $\rm 400-4000\ M_{\odot}$. The studies among others have posited that the VMS that forms in this way could be a potentially evolve into an IMBH (Freitag2007; Freitag2008; Goswami2012; Gonza2024).
@@
Simulations of the evolution star cluster models carried out using the Monte Carlo MOCCA code (Hypki2013; Giersz2013) have shown that massive clusters with high central densities ($\rm\rho_{c}\gtrsim 10^{7}\ \mathrm{M}_{\odot}\ pc^{-3}$) are likely to form a VMS through the runaway collisions of MS stars (Giersz2015; Askar2017; Hong2020; Maliszewski2022; Askar2023). In order to trigger this pathway for VMS formation and subsequently IMBH formation, mass segregation of the most massive stars in the cluster must occur on a timescale shorter than their lifetime which is about 3 Myr (Gurkan2004). High central densities and large sizes of MS stars increases their likelihood to collide and form a VMS. From thousands of star cluster models simulated as part of the MOCCA-Survey Database I (Askar2017a), it has been shown that a rapid collisional runaway leading to IMBH formation always occurs when the initial central density exceeds $\rm\rho_{c}\gtrsim 10^{7}\ \mathrm{M}_{\odot}\ pc^{-3}$ (see Fig. 1 in Hong2020). The mass of the VMS that forms through this pathway is of the order $10^{3}-10^{4}\ \mathrm{M}_{\odot}$ and depends on the total cluster mass and initial central density. Although the formation of a VMS and subsequently an IMBH through a rapid collisional runaway has been extensively explored in previous MOCCA simulations of massive and dense star clusters, these results had not been directly verified with N-body simulations for clusters with extremely high central densities until now.
@@
In addition to employing the Henon1971 Monte Carlo method to treat two-body relaxation, MOCCA utilizes the fewbody code, a direct N-body integrator for small-N gravitational dynamics, to determine the outcomes of close binary-single and binary-binary interactions (Fregeau2004). For physical collisions between two MS stars in MOCCA, a sticky-sphere assumption with no mass loss is applied. Similar to NBODY6++GPU, the MOCCA code models stellar and binary evolution using prescriptions from the SSE/BSE rapid population synthesis codes (Hurley2000; Hurley2002), which have been subsequently improved and extended with more recent developments (see Kamlah2022a, and references therein). While the MOCCA code can model a galactic tidal field using a point-mass approximation, no external tidal field was applied in the simulations carried out in this study, in order to remain consistent with the direct N-body run. A detailed description of the latest MOCCA features is provided in Hypki2022. For this work, MOCCA was updated to incorporate the improved treatment of stellar radii and the rejuvenation process for a VMS that have been described in Appendices A, B and C. Section 3.2 outlines the initial conditions and code modifications applied to MOCCA, while Section 4.3 presents the results on VMS growth and IMBH formation in the MOCCA simulation.
@@ 3.1 Star cluster parameters
Half-mass radius ($r_{\rm hm}$) $0.10\penalty 10000\ \mathrm{pc}$
Half-mass average density ($\rho_{\rm hm}$) $6.99\cdot 10^{7}\penalty 10000\ \rm M_{\odot}/{\rm pc}^{3}$
Density profile King ($W_{0}$) 66 (1)
@@
IMF, min/max stellar mass $0.08-150\penalty 10000\ \rm M_{\odot}$ (2)
Initial cluster mass ($M_{\rm i}$) $5.86\cdot 10^{5}\penalty 10000\ \rm M_{\odot}$
@@
In this paper we present one selected star cluster model, all relevant global parameters are given in Table 1. The dynamical evolution of this system is followed by Nbody6++GPU as well as MOCCA, for the purpose of comparison. The initial model was generated by Mcluster (Kuepper2011; Leveque2021). The initial number of stars is set to $10^{6}$, with no initial (primordial) binaries. While such binaries in principle can be included (see e.g. Wang2016; Rizzuto2021) they are computationally very expensive and also many free parameters for their initial configuration are added. We postpone models with many primordial binaries and a more extended parameter study of initially highly compact star cluster models to future work; in our simulations, however, binaries form and evolve dynamically.The cluster density is modelled via a King1962 profile with central dimensionless well $W_{0}=6$. Initial stellar masses are distributed according to a Kroupa2001 initial mass function limited between $0.08-150\penalty 10000\ \rm M_{\odot}$. All details about our cluster model are summarised in Table 1. The cluster model is initially in virial equilibrium, with no initial mass-segregation, rotation , or any level of fractality (Goodwin2004). We do not use an external tidal field i.e., an external gravitational field from the host galaxy, since we are more interested in the dynamical evolution near and inside the half-mass radius. Still, it is possible that stars escape, if their total energy is positive and their distance from the cluster center is larger than 20 initial half-mass radii Aarseth2003a.
Crossing time and relaxation time taken at the half-mass radius of a star cluster ($t_{\rm cr,hm}$ $t_{\rm rx,hm}$) are important parameters determining the general evolution of a star cluster (Binney1987; Binney2008). The crossing time corresponds to the time in which a star with a typical velocity travels through the cluster under the assumption of virial equilibrium. The half-mass relaxation time corresponds to the time over which the cumulative effect of stellar encounters on the velocity of a star becomes comparable to the star’s initial velocity. In our star cluster model the initial relaxation time is 7.017.01\, Myr, and the initial crossing time is $5.9\cdot\ 10^{-4}\,$ Myr.
The mass segregation timescale is important to describe the initial segregation of our star cluster model (Kroupa1995). The mass segregation timescale is defined as:
@@
where $\overline{m}$ is the initial average mass of the stars in the cluster, while $m_{\rm max}$ is the largest star mass in the cluster, which are 0.58 $\penalty 10000\ \rm M_{\odot}$ and 148.49 $\penalty 10000\ \rm M_{\odot}$, respectively. This, implies a segregation time of $t_{\rm ms}=0.03$ Myr. Moreover, given the large density of the cluster model, this means that encounters between large masses occurs relatively early in the star cluster dynamical history.
Another relevant timescale to describe the stellar dynamics is the (average) collision time, which quantifies the rate of collisions between stars in the evolution cluster (Binney1987; Binney2008) is $t_{\rm coll}=5.35\cdot\ 10^{-3}\,$ Myr. According to the scenario proposed by Escala2021, from the comparison of the relaxation and collision times to the age of the system $\tau$, it is possible to determine if collisions will play an important role in building a massive central object and the long term stability of the cluster. To test this, Vergara2023 performed idealized (equal mass) numerical simulations and in order to quantify this transition, defined a critical mass for $t_{\rm coll}=\tau$, which can be expressed as follows:
@@
where the effective cross-section is given by $\Sigma_{0}=16\sqrt{\pi}R_{*}^{2}(1+\Theta)$, the Safronov number is defined as $\Theta=9.54\left(M_{*}R_{\odot}/\penalty 10000\ \rm M_{\odot}R_{*}\right)\left(100\penalty 10000\ \rm{km\,s^{-1}}/\sigma\right)^{2}$, and $n$ is the numeric density. Vergara2023 found that the fraction of the total mass in the central object dramatically increase for $t_{\rm coll}<\tau$, being up to 50% for the most extreme runs.
Fig. 1, shows the initial mass of the cluster against the half-mass radius and includes a color bar with the average logarithm of the number of stars. It also displays the relaxation time and the collision time (or critical mass), when both are equals to a time evolution (i.e. age) of $\tau=10\,$ Gyr, which we adopt here as an upper limit for the evolutionary timescales in the simulations we are comparing with. We compare the respective simulations for several grids of direct $N$-body shown as circle symbols in plot representing initial conditions from P+99: PortegiesZwart1999, F+13: Fujii2013, K+15: Katz2015, W+15: Wang2015, M+16: Mapelli2016, S+17: Sakurai2017, R+18: Reinoso2018, P+19: Panamarev2019, R+20: Reinoso2020, D+20: DiCarlo2020a, R+20: Rastello2021, R+21: Rizzuto2021, B+21: Banerjee2021, G+21: Gieles2021, V+21: Vergara2021, K+22: Kamlah2022b, R+23: Rizzuto2023, V+23: Vergara2023, and AS+23: ArcaSedda2023a; ArcaSedda2024a; ArcaSedda2024, Monte-Carlo shown as square symbols in plot representing initial conditions from A+17: Askar2017a, R+19: Rodriguez2019, K+20: Kremer2020a, M+22: Maliszewski2022, and R+22: Rodriguez2022 and hybrid simulations diamond symbols in plot representing initial conditions from W+21: Wang2021, W+22: Wang2022a, and W24: Wang2024. The latter are classified as hybrid $N$-body simulations, because they combine particle-particle (e.g. direct $N$-body) and particle-tree (e.g. Barnes-Hut tree) methods to speed-up calculations, and exploit the advantage of a highly parallelised implementation using the code PeTaR (Wang2020a).